Dont get this at all -
The variable Z varies jointly with x and y. Use the given values to write an equation relating x, y, and z. Then find Z when x = -4 and y = 7.
The equation would be z = kxy - right???
Now
x = 1, y = 1/3, z = 5
x = -6, y = 3, z = 2/5
Dont know what to do at all
HELP

Answer by stanbon(75887)

The variable Z varies jointly with x and y.
Use the given values to write an equation relating x, y, and z.
Then find Z when x = -4 and y = 7.
The equation would be z = kxy - right???
That is correct
-----------
If x=-4 and y = 7, then z = k(-4)(7); z=-28k
------------------
Now
x = 1, y = 1/3, z = 5
5 = k(1*(1/3)), k= 15
---------
x = -6, y = 3, z = 2/5
(2/5) = k(-6)(3)
(2/5) = -18k
(1/5) = -9k
k= -1/45
============
Cheers,
Stan H.